The amplitude of a spring-mass system in simple harmonic motion is 8 m. Calculate the value of
displacement x where the magnitude of the velocity is 9/16 of its maximum value.
1
Expert's answer
2020-05-19T10:49:52-0400
"E=0.5kA^2=0.5mV^2=0.5kx^2+0.5mv^2"

"0.5mv^2=\frac{9^2}{16^2}E=0.5kA^2\frac{9^2}{16^2}"

"A^2=x^2+A^2\frac{9^2}{16^2}\to 8^2=x^2+8^2\frac{9^2}{16^2}"

"x=6.6\ m"